OpenStudy (anonymous):

name the property illustrated x+(3+2x)=x+(2x+3)

OpenStudy (anonymous):

maybe this will help? http://www.algebrahelp.com/

OpenStudy (anonymous):

Associative Property An operation is associative if you can group numbers in any way without changing the answer. It doesn't matter how you combine them, the answer will always be the same. Addition and multiplication are both associative. Here are some addition examples: 1 + (2 + 3) = (1 + 2) + 3 1 + (5) = (3) + 3 6 = 6. (-1 + 66) + 14 = -1 + (66 + 14) (65) + 14 = -1 + (80) 79 = 79. More generally, a + (b + c) = (a + b) + c. Here is a multiplication example: 2 * (4 * 3) = (2 * 4) * 3 2 * (12) = (8) * 3 24 = 24. In general terms, that's a * (b * c) = (a * b) * c.

OpenStudy (anonymous):

It looks like the associative property of addition and also the commutative property of addition

OpenStudy (anonymous):

Actually nix that, it's just the commutative property.

OpenStudy (anonymous):

i think it's both Commutative Property and associative An operation is commutative if you can change the order of the numbers involved without changing the result. Addition and multiplication are both commutative. Subtraction is not commutative: 2 - 1 is not equal to 1 - 2. Here are some examples of the commutative properties of addition and multiplication: 88 + 65 = 65 + 88 153 = 153. 12 * 13 = 13 * 12 156 = 156. More generally, a + b = b + a, and a * b = b * a. From the Math Forum:

OpenStudy (anonymous):

In the problem the parenthesis never change what they are around (the 2x and the 3) so I think this is just the commutative property of addition.